According to the data reported by the New York State Department of Health regarding West Nile Virus for the years 2000-2004, the least squares line equation for the number of reported dead birds (x) versus the number of human West Nile virus cases (y) is 

 = −10.2638 + 0.0491x.

 If the number of dead birds reported in a year is 554, how many human cases of West Nile virus can be expected? (Round your answer to the nearest whole number.)
Answer #1

This is a simple case of finding the predicted value of \(y\) that is the predicted value of the number of human cases of West Nile Virus that can be expected dependent on our least square model.

\(\hat{y}=-10.2638+0.0491 x\)

Put \(x=554\) in the above equation to get:

\(\hat{y}=-10.2638+0.0491 * 554\)

\(\hat{y}=-10.2638+27.2014=16.94\)

Which is 17 rounded to the nearest whole number.

Therefore the number of human cases of West Nile that can be expected are 17